3-11. At some instant the velocity components of an electron moving between two charged parallel
plates are vx = 1.5105 m/s and vy = 0.30104 m/s. If the electric field between the plates is given by
E = j1.2104 N/C, (a) what is the acceleration of the electron? (b) When the x-coordinate of the
electron has changed by 2.0 cm what will be the velocity of the electron?
1
Expert's answer
2019-12-18T11:26:12-0500

We are given

Vx= 1.5*105 m/s Vy=0.3*104 m/s

a) As electric field is in the positive j then the force acting on the electron will be in negative j direction

F=eE

ma=eE

a=eE/m

a=1.6*10-19*1.2*104/(9.1*10-31)

a=2.1*1015 m/s2 in negative j

b) as in positive x axis there is no electric field so no force will act on the electron

so time taken by the electron to travel 2 cm is 2*10-3/(1.5*105) = 1.8*10-8sec

velocity in the x direction will remain to be the same

velocity in y= uy-at

=0.3*104 - 2.1*1015*1.8*10-8

=0.3*104 - 3.78*107

=3.78*107 in negative j